a sheet of steel is to be cut into panels. The total area of the sheet is 5m2. Panel C is 4 m2. Panels A and B are identical in size.
using algebra, find the area of panel A. show all working out

Answer by mananth(16946)

let panel A be x square meter
panel C is 4 sq. meter
Panel B is x square meter
total area of sheet is 5 sq. meter
x+x+4 =5
2x+4 =5
2x=1
x=1/2
Panel A is 1/2 square meter